OpenStudy (anonymous):

Need help in differentiation. Differentiate w.r.t. x. 2[sqrt(x)+2]^(1/2)

OpenStudy (anonymous):

In an easier-to-understand form - \[2(\sqrt{x}+2)^.5\]

OpenStudy (anonymous):

idk:(

OpenStudy (anonymous):

f(x)=2[sqrt(x)+2]^(1/2) f '(x)=2(1/2)(sqrtx+2)^-(1/2) D(sqrtx +2) =[1/(sqrtx +2)^(1/2)] (1/2)x^-(1/2) =[1/(sqrtx +2)^(1/2)] (1/2)x^-(1/2) =1/[sqrtx (sqrtx +2)^.5]

OpenStudy (anonymous):

\[\Large \begin{array}{l} 2\sqrt {\sqrt x + 2} \\ = 2\sqrt {{x^{1/2}} + 2} \\ = 2{\left( {{x^{1/2}} + 2} \right)^{1/2}}\\ f(x) = 2{(x)^{1/2}}\\ f'(x) = {x^{ - 1/2}}\\ g(x) = \sqrt x + 2 = {x^{1/2}} + 2\\ g'(x) = \frac{1}{2}{x^{ - 1/2}}\\ \frac{d}{{dx}} = {\left( {{x^{1/2}} + 2} \right)^{ - 1/2}} \cdot \frac{1}{2}{x^{ - 1/2}}\\ = \frac{1}{2} \cdot \frac{1}{{\sqrt {{x^{1/2}} + 2} }} \cdot \frac{1}{{\sqrt x }}\\ = \frac{1}{{2\sqrt {\sqrt x + 2} \cdot \sqrt x }} \end{array}\]

OpenStudy (anonymous):

f(x)=2[sqrt(x)+2]^(1/2) f '(x)=2(1/2)(sqrtx+2)^-(1/2) D(sqrtx +2) =[1/(sqrtx +2)^(1/2)] (1/2)x^-(1/2) =[1/(sqrtx +2)^(1/2)] (1/2)x^-(1/2) =1/[2sqrtx (sqrtx +2)^.5]